OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a=7 b=8 c=7

OpenStudy (anonymous):

use cosine laws...

OpenStudy (anonymous):

ok, so far i know to use c^2=a^2+b^2-2abcosC. can u see if i'm doing this right?

OpenStudy (anonymous):

yes y not..show your work..

OpenStudy (anonymous):

ok, I have 7^2=4^2+8^2-2*4*8cosC => 49-16-64=64cosC

OpenStudy (anonymous):

use correct values..you plug in a=4 but it is a=7

OpenStudy (anonymous):

oh ok, so if a=7, c=4?

OpenStudy (anonymous):

still you plug in wrong no if a=7 and c=4 but it is true if a=4 andc =7

OpenStudy (anonymous):

oh my bad, i meant to put a=4, b=8, and c=7

OpenStudy (anonymous):

so 7^2=4^2+8^2-2*4*8cosC => 49-16-64= 64cosC

OpenStudy (anonymous):

xactly right plugging this time..

OpenStudy (anonymous):

thanks. and to find 64cosC, do i replace it with 180?

OpenStudy (anonymous):

no...now simplyfy the left side and then divide by 64 to free cosC..then take cos^-1

OpenStudy (anonymous):

ok, so 49-16-64=-31, (-31/64)= -.48437(or -0.5) , cos^-1(-0.5)=120

OpenStudy (anonymous):

i got it now. I actually gotta use the formula b^2=a^2+c^2-2accosB. so i have 8^2=4^2+7^2-2*4*7cosB => -1/-56= cos^-1(.0178)=88.9=89 degrees.

OpenStudy (anonymous):

thank u. i think i can do the rest of it on my own now.

OpenStudy (anonymous):

all i can do is use the sine laws to solve the rest.

OpenStudy (anonymous):

okay..its allright..